Bill has decided to begin investing for retirement. He is going to put his tax raturn of $7,328 into the account today and then set up monthly contributions from his paycheck in the amount of $385. The money will go into an account that earns 7.7 percent annual interest (compounded monthly). How much will he have in the account after 20 years? Show using excell cell functions

Nper
Rate
PV
FV
PMT
Answer #1

Nper = 20*12 = 240

Rate = 7.7%/12

PV = -7328

PMT = -385

Solve for FV

=FV(7.7%/12, 240, -385, -7328) = $252516.23